# JWST detection of a carbon dioxide dominated gas coma surrounding interstellar object 3I/ATLAS

### v1

@@ 5 Discussion
Our JWST NIRSpec observations reveal that 3I/ATLAS containins a substantial volatile ice inventory, with a gas coma unusually rich in CO2 relative to H2O and CO. In the absence of clear detections of other gases, it is reasonable to infer that CO2 outgassing provides the dominant driving force for 3I’s nucleus activity, and is responsible for launching dust grains away from the nucleus to produce the distinctive scattered light coma observed at m and shorter wavelengths.
@@
The interstellar object 3I/ATLAS has a coma CO2/H2O ratio of , which is 16 times larger than expected for its heliocentric distance, based on the fit to previously observed cometary data (Figure 6). This corresponds to away from the trend line, and shows that 3I’s coma CO2/H2O ratio is unusually high. The only other comet known to have a CO2/H2O ratio so far outside the normal Solar System trend is C/2016 R2 (PanSTARRS) (A. J. McKay et al., 2019). C/2016 R2 is considered to be one of the most peculiar comets ever observed, as a result of its large hypervolatile content (N. Biver et al., 2018; M. A. Cordiner et al., 2022) and correspondingly high CO/H2O ratio. The CO/H2O ratio of in 3I/ATLAS, on the other hand, is more compatible with previous cometary observations, which have values between –4 au (O. Harrington Pinto et al., 2022). Intriguingly, our CO/H2O ratio is within the range of values (1.3–1.6) measured in 2I/Borisov at au, although the H2O production rate was observed to be falling rapidly around the time of those observations (D. Bodewits et al., 2020).
@@
Due to the presence of water ice in cometary comae at 3 au (e.g. E. Lellouch et al., 1998; H. Kawakita et al., 2004; S. Protopapa et al., 2018), previous measurements of the gas CO2/H2O ratio using larger spectroscopic apertures (e.g. T. Ootsubo et al., 2012) could have been impacted by icy grain sublimation, thus reducing the observed ratio. It is difficult to assess the full impact this would have on the points around –4 au in Figure 6. We therefore recommend further observations of the volatile content of 3I/ATLAS during its post-perihelion observing window near au, to facilitate the study of its bulk ice composition.
We used the model of J. J. Cowan & M. F. A’Hearn (1979) to calculate the CO2, H2O, and CO active surface fractions, assuming an infrared emissivity of 0.95, albedo of 4% and nucleus radius of 2.8 km (D. Jewitt et al., 2025b). The Small-Bodies-Node/ice-sublimation code (M. Van Selous & M. Kelley, 2021) was used to calculate the average sublimation rate per unit area, , at = 3.32 au. The active area is derived by dividing our terminal values by , and the nucleus active fractional area is found by dividing the active area by the nucleus surface area. The lower limits on active fractional area for CO2, CO, and H2O are , , and , respectively. According to the analysis of D. Jewitt et al. (2025b), our measured CO2 active area of 2.7 km2 is sufficient to drive the development of 3I’s observed dust coma, even in the presence of large ( m) dust grains.
@@
A lower active area for H2O than CO2 could be partly explained by the higher sublimation temperature of H2O, if the internal temperature of most of the nucleus () was in the range (but still close enough to to allow some H2O sublimation), at the time of our observations. Sublimation of H2O may therefore become more fully activated as 3I/ATLAS moves closer to the Sun (T. H. Puzia et al., 2025), in which case a more accurate picture of the nucleus composition will be obtained. The very high CO2/H2O ratio observed by JWST could therefore indicate that is lower than that experienced by typical Solar System comets at a similar . This could arise as a result of a higher albedo or lower thermal conductivity of the nucleus surface layer compared with typical comets, leading to reduced heating or heat penetration. Higher albedo could be caused by a more ice-rich surface composition than normal, whereas lower thermal conductivity could arise from the presence of a volatile-depleted crust/mantle (A. Guilbert-Lepoutre et al., 2015). The latter was hypothesized for 1I/‘Oumuamua, as a result of irradiation by cosmic rays during the object’s interstellar passage (A. Fitzsimmons et al., 2018). Accounting for the lower volatility of H2O using the (J. J. Cowan & M. F. A’Hearn, 1979) ice sublimation model, we predict CO2/H2O at au, which is still an order of magnitude larger than other comets observed near au (Figure 6).
@@ 6 Conclusion
We performed JWST NIRSpec imaging spectroscopy of the interstellar object 3I/ATLAS at au on the inbound leg of its flight through the Solar System. Rovibrational emission bands were detected of CO2, CO and H2O, in addition to dust and ice solid-state features, demonstrating the presence of a substantial, gas- and ice-rich coma comparable to that of comets from our own Solar System. The CO2 band at 4.3 m was particularly strong. Gas production rates were derived using PSG spectral models, resulting in a CO2/H2O mixing ratio of , which is above the trend as a function of observed in long-period and Jupiter-family comets (excluding the peculiar C/2016 R2), and suggests the possibility of an intrinsically CO2-rich nucleus. Such a high CO2/H2O ratio has never before been observed in a comet between –4 au. The combined capabilities of the JWST and Vera C. Rubin Legacy Survey of Space and Time (Ž. Ivezić et al., 2019) will facilitate additional observations of Solar System comets at such distances, to help improve the statistics and confirm whether 3I/ATLAS is as unusual as it appears.
@@ Appendix A Radially-Enhanced Flux Maps
Figure 7 shows “ enhanced” versions of the flux maps in Figure 2, where the dilution of the observed column density due to quasi-spherical expansion of the coma (which is proportional to , where is the sky-projected distance from the nucleus), has been divided out.
Appendix B Retrieved Gas Rotational Temperature Curves
Rotational temperatures as a function of nucleocentric distance, retrieved from independent spectral line modeling of each observed species (Section 4) are shown in Figure 8.
Appendix C 13CO2 in 3I/ATLAS
@@ Appendix B Retrieved Gas Rotational Temperature Curves
Examination of Figure 1 indicates the presence of 13CO2 alongside 12CO2, affording the first opportunity to test the 12CO2/13CO2 ratio in an interstellar object. Compared to the much stronger 12CO2 emission, the 13CO2 band is only clearly detected in the the nucleus-centered aperture. Unfortunately, the 12CO2 in this region suffers from optical depth effects, and the low spectral resolution introduces additional difficulties in disentangling the contributions from each isotopologue. We therefore focused on analyzing the spectra from the first annular sector, where optical depth effects are reduced. We used the PSG to retrieve production rates s-1 and s-1. This corresponds to a 12C/13C lower limit of 77 (3), which is formally consistent with the terrestrial value of 89. Additional observations of 13CO2 in 3I/ATLAS at higher spectral resolution and signal-to-noise will be invaluable in separating it from its optically thick 12CO2 counterpart, in order to place improved constraints on the CO2 isotopic ratio.

### v3

@@ 5 Discussion
Our JWST NIRSpec observations reveal that 3I/ATLAS contains a substantial volatile ice inventory, with a gas coma unusually rich in CO2 relative to H2O and CO. In the absence of clear detections of other gases, it is reasonable to infer that CO2 outgassing provides the dominant driving force for 3I’s nucleus activity, and is responsible for launching dust grains away from the nucleus to produce the distinctive scattered light coma observed at m and shorter wavelengths.
@@
The interstellar object 3I/ATLAS has a coma CO2/H2O ratio of , which is 18 times larger than expected for its heliocentric distance, based on the fit to previously observed cometary data (Figure 6). This corresponds to away from the trend line, and shows that 3I’s coma CO2/H2O ratio is unusually high. The only other comet known to have a CO2/H2O ratio so far outside the normal Solar System trend is C/2016 R2 (PanSTARRS) (A. J. McKay et al., 2019). C/2016 R2 is considered to be one of the most peculiar comets ever observed, as a result of its large hypervolatile content (N. Biver et al., 2018; M. A. Cordiner et al., 2022) and correspondingly high CO/H2O ratio. The CO/H2O ratio of in 3I/ATLAS, on the other hand, is more compatible with previous cometary observations, which have values between –4 au (O. Harrington Pinto et al., 2022). Intriguingly, our CO/H2O ratio is within the range of values (1.3–1.6) measured in 2I/Borisov at au, although the H2O production rate was observed to be falling rapidly around the time of those observations (D. Bodewits et al., 2020).
@@
Due to the presence of water ice in cometary comae at 3 au (e.g. E. Lellouch et al., 1998; H. Kawakita et al., 2004; S. Protopapa et al., 2018), previous measurements of the gas CO2/H2O ratio using larger spectroscopic apertures (e.g. T. Ootsubo et al., 2012) could have been impacted by icy grain sublimation, thus reducing the observed ratio. It is difficult to assess the full impact this would have on the points around –4 au in Figure 6, so further observations of 3I/ATLAS are recommended, closer to perihelion.
@@
The active sublimating surface area for each of our detected gases is calculated in Appendix C. Following the analysis of D. Jewitt et al. (2025), the CO2 active area of 3.1 km2 is sufficient to drive the development of 3I’s observed dust coma, even if the dust grains are relatively large ( m) in size. The relatively small active area for H2O (2.2 km2) could be partly explained by the relatively high sublimation temperature of H2O (M. Womack et al., 2017), if the internal temperature of most of the nucleus () was in the range (but still close enough to to allow some H2O sublimation), at the time of our observations. Sublimation of H2O may therefore become more fully activated as 3I/ATLAS moves closer to the Sun (T. H. Puzia et al., 2025), in which case a more accurate picture of the nucleus composition will be obtained. The very high CO2/H2O ratio observed by JWST could therefore indicate that is lower than that experienced by typical Solar System comets at a similar . This could arise as a result of a higher albedo or lower thermal conductivity of the nucleus surface layer compared with typical comets, leading to reduced heating or heat penetration. Higher albedo could be caused by a more ice-rich surface composition than normal, whereas lower thermal conductivity could arise from the presence of a volatile-depleted crust/mantle (A. Guilbert-Lepoutre et al., 2015). The latter was hypothesized for 1I/‘Oumuamua, as a result of irradiation by cosmic rays during the object’s interstellar passage (A. Fitzsimmons et al., 2018). Accounting for the lower volatility of H2O using the (J. J. Cowan & M. F. A’Hearn, 1979) ice sublimation model, we predict CO2/H2O at au, which is still an order of magnitude larger than other comets observed near au (Figure 6).
@@ 6 Conclusion
We performed JWST NIRSpec imaging spectroscopy of the interstellar object 3I/ATLAS at au on the inbound leg of its flight through the Solar System. Rovibrational emission bands were detected of CO2, CO and H2O, in addition to dust and ice solid-state features, demonstrating the presence of a substantial, gas- and ice-rich coma comparable to that of comets from our own Solar System. The CO2 band at 4.3 m was particularly strong. The CO2/H2O mixing ratio of is above the trend as a function of observed in long-period and Jupiter-family comets (excluding the peculiar C/2016 R2), and suggests the possibility of an intrinsically CO2-rich nucleus. Such a high CO2/H2O ratio has never before been observed in a comet between –4 au. The combined capabilities of the JWST and Vera C. Rubin Legacy Survey of Space and Time (Ž. Ivezić et al., 2019) will facilitate additional observations of Solar System comets at such distances, to help improve the statistics and confirm whether 3I/ATLAS is as unusual as it appears.
@@ Appendix A Radially-Enhanced Flux Maps
Figure 7 shows “ enhanced” versions of the flux maps in Figure 2, where the dilution of the observed column density due to quasi-spherical expansion of the coma (which is proportional to , where is the sky-projected distance from the center of the nucleus pixel), has been divided out. The average value of was calculated within each pixel using a -point cartesian sub-sampling.
Appendix B Spectral Modeling
Gas production rates () and rotational temperatures (), were derived as a function of distance from the nucleus for CO2, CO, and H2O, using optimal estimation routines as part of the Planetary Spectrum Generator (PSG; G. L. Villanueva et al. 2018, based on synthetic fluorescence models described by G. Villanueva et al. 2025). We performed baseline fitting to subtract the (nucleus + dust) continuum based on fits to the spectral regions immediately adjacent to the lines of interest (Figure 4), adopting the conservative strategy of obtaining an good fit within the noise, using a minimum number of free parameters. A 3rd-order polynomial was found to be sufficient for describing the continua underneath and surrounding the CO2 and CO features, whereas for H2O, we tried various analytic functions to produce a good fit, including the wing of the 3 m ice band. Although a 5th order polynomial was found to be sufficient to reproduce the shape of the continuum in this region, a lack of formal constraints across the 2.6–2.8 m region spanning the H2O 2.7 m band led to increased uncertainties on . Therefore, for H2O, we adopted a more physically constrained continuum model, formed as the product of a linear slope and a sigmoid function:
where is the linear continuum normalization point, is the linear slope, is the y-intercept point, is the depth of the 3-µm band, controls the slope of the band edge, and controls the wavelength of the band edge. The (variable) exponential onset of the sigmoid function effectively matches the quasi-Gaussian shape of the blue wing of the 3 m H2O ice band (G. Leto & G. A. Baratta, 2003). By simultaneously optimizing the continuum and spectral line models, uncertainties in the fitted continuum shapes were included in the uncertainties derived for our best-fitting production rates. After experimenting with alternative functional forms for the continuum in the vicinity of our observed spectral lines, we found that the derived production rates remained consistent, within the errors.
The CO2 and CO spectral regions were modeled using the methods described by C. E. Woodward et al. (2025). The gas outflow velocity was set at 0.44 km s-1, based on the standard relationship (e.g. T. Ootsubo et al., 2012). Molecular photolysis rates appropriate for the active Sun were incorporated from W. F. Huebner & J. Mukherjee (2015). Pixels close to the nucleus can be affected by significant line opacity and PSF-related flux losses, which are difficult to accurately model. Therefore, after modeling the average spectrum within the nucleus-centered aperture, we proceeded with a “-curve” analysis, deriving the production rates and rotational temperatures as a function of within successive (independent) partial annular sectors centered on the brightest (nucleus-containing) pixel. To avoid flux losses from pixels at the very edge of the IFU, and to focus on a uniform coma angular region, a inscribed angle was used for all partial annuli, with radial bounds parallel to the north-east and south-east edges of the IFU; see inset diagram in Figure 5. We found a annulus width to provide sufficient signal-to-noise for all species. Given the very high SNR for CO2, we also generated a higher-resolution curve for this species using a annulus width. The resulting production rates and rotational temperatures as a function of are shown in Figure 5.
For the m region, H2O spectral models were generated using PSG, while the continuum was determined using Equation B1, combined with the =5000 Solar spectral model of R. C. Bohlin et al. (2014). Additional (weak) spectral contributions from coma CO2 to this region were added based on our best fit to the 4.3 m band. The H2O production rate and rotational temperature were retrieved using a Gaussian processes approach within the George software package (S. Ambikasaran et al., 2015), and uncertainties were derived using a Markov Chain Monte Carlo algorithm (emcee; D. Foreman-Mackey et al., 2013). The retrieved rotational temperatures as a function of nucleocentric distance for each observed species are shown in Figure 8.
@@ Appendix B Spectral Modeling
Examination of Figure 1 indicates the presence of 13CO2 alongside 12CO2, affording the first opportunity to test the 12CO2/13CO2 ratio in an interstellar object. Compared to the much stronger 12CO2 emission, the 13CO2 band is only clearly detected in the the nucleus-centered aperture. Unfortunately, the 12CO2 in this region suffers from optical depth effects, and the low spectral resolution introduces additional difficulties in disentangling the contributions from each isotopologue. We therefore focused on analyzing the spectra from the first annular sector, where optical depth effects are reduced. We used the PSG to retrieve production rates s-1 and s-1. This corresponds to a 12C/13C lower limit of 63 (3), which is formally consistent with the terrestrial value of 89. Additional observations of 13CO2 in 3I/ATLAS at higher spectral resolution and signal-to-noise will be invaluable in separating it from its optically thick 12CO2 counterpart, in order to place improved constraints on the CO2 isotopic ratio.
Appendix C Active Sublimating Surface Areas
We used the cometary ice thermal sublimation model of J. J. Cowan & M. F. A’Hearn (1979) to calculate the CO2, H2O, and CO active surface areas, assuming an infrared emissivity of 0.95, albedo of 5% and nucleus radius of km (D. Jewitt et al., 2025). The Small-Bodies-Node/ice-sublimation code (M. Van Selous & M. Kelley, 2021) was used to calculate the average sublimation rate per unit area, , at = 3.32 au, assuming a non-rotating, spherical nucleus. The active area is derived by dividing our terminal values by , resulting in values of 3.1 km2 for CO2, 2.2 km2 for H2O, and 0.2 km2 for CO. The active fractional area of the nucleus for each species is found by dividing these active areas by the (assumed) nucleus surface area. The corresponding lower limits on the active fractional area for each ice are , , and , respectively.